OpenStudy (anonymous):

Sorry, I'm in 8th grade, just starting Algebra 1. I need a refresher. Solve for k, 4/5k + 7/10 = 13/15k - 3/5

OpenStudy (anonymous):

make your life easy by multiplying both sides by 30 to get rid of the fractions

OpenStudy (anonymous):

\[\frac{4}{5}k + \frac{7}{10} = \frac{13}{15}k -\frac{ 3}{5}\]

OpenStudy (anonymous):